# SeisFusion: Constrained Diffusion Model with Input Guidance for 3D Seismic Data Interpolation and Reconstruction

### v1

@@ Abstract
Geographical, physical, or economic constraints often result in missing traces within seismic data, making the reconstruction of complete seismic data a crucial step in seismic data processing. Traditional methods for seismic data reconstruction require the selection of multiple empirical parameters and struggle to handle large-scale continuous missing data. With the development of deep learning, various neural networks have demonstrated powerful reconstruction capabilities. However, these convolutional neural networks represent a point-to-point reconstruction approach that may not cover the entire distribution of the dataset. Consequently, when dealing with seismic data featuring complex missing patterns, such networks may experience varying degrees of performance degradation. In response to this challenge, we propose a novel diffusion model reconstruction framework tailored for 3D seismic data. To constrain the results generated by the diffusion model, we introduce conditional supervision constraints into the diffusion model, constraining the generated data of the diffusion model based on the input data to be reconstructed. We introduce a 3D neural network architecture into the diffusion model, successfully extending the 2D diffusion model to 3D space. Additionally, we refine the model’s generation process by incorporating missing data into the generation process, resulting in reconstructions with higher consistency. Through ablation studies determining optimal parameter values, our method exhibits superior reconstruction accuracy when applied to both field datasets and synthetic datasets, effectively addressing a wide range of complex missing patterns. Our implementation is available at https://github.com/WAL-l/SeisFusion.
@@ I Introduction
Traditional methods, although theoretically capable of data interpolation and reconstruction[17], heavily depend on manual parameter selection and struggle with interpolating large continuous missing data[18]. In recent years, with the development of deep learning in recent years, the use of deep learning for seismic data processing has attracted a lot of attention [19] [20] [21], and neural networks with different structures have been proposed for the reconstruction of seismic data[22, 23]. In the realm of 2D reconstruction, Wang et al. [14] advocated for using CAE to interpolate missing seismic data, while Chai et al. [24] employed the UNet network for seismic data reconstruction. However, since real seismic data is inherently 3D, focusing solely on inline slices for 2D reconstruction neglects crucial crossline information. One by one reconstruction of 2D slices can introduce discontinuities in stacked slices, particularly in datasets featuring large-scale or continuous missing data and other complex scenarios. Therefore, opting for 3D reconstruction is a more pragmatic approach. Chai et al.[25] utilized an end-to-end 3D UNet for data reconstruction, observing superior performance compared to 2D methods. Qian et al.[26] proposed a deep tensor autoencoder, introducing tensor backpropagation, which exhibited commendable performance in reconstructing irregular data. As GAN[27] showcase potent generative capabilities, utilizing improved GAN networks for seismic data interpolation has gained traction. Dou et al.[16] introduced a multidimensional adversarial generative adversarial network (MDA GAN), incorporating three discriminators and refining the loss function, resulting in robust reconstruction performance. Yu et al.[28] advocated for a CWGAN for data reconstruction, with results demonstrating its superiority over 3D UNet.
@@
Figure 1: The diffusion model, where $q$ represents the forward encoding training process, and $p$ represents the reverse sampling generation process.
@@
1. Introduced conditional supervision constraints into the diffusion model, constraining the generated data of the diffusion model based on input conditions.
@@
3. Improved the model’s generation and reconstruction process by guiding the model to generate reconstructed data through inputting the missing data to be reconstructed into the generation and reconstruction process, resulting in the generation of reconstructed data with higher consistency.
@@ II-A Forward Encoding Training Process:
$q(x_{t}\mid x_{t-1})=\mathcal{N}(\mu_{t},\beta_{t}\mathbf{I})$ (1)
Where $\beta_{t}$ is obtained from the existing variance table, typically set to linearly increase from 0.0001 to 0.002. $\mu_{t}$ is the mean, according to Nichol et al.[36], $\mu_{t}=\sqrt{1-\beta_{t}}x_{t-1}$. Therefore, Equation (1) can be rewritten as:
@@
At the same time, $x_{t}$ can be obtained using reparameterization [46].
$x_{t}=\sqrt{\overline{\alpha}_{t}}x_{0}+\epsilon_{t}\sqrt{(1-\overline{\alpha}_{t})},\epsilon_{t}\sim\mathcal{N}(0,\mathbf{I})$ (4)
where $\alpha_{t}=1-\beta_{t},\overline{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$. $\epsilon_{t}$ is sampled from a Gaussian distribution.
@@
According to Ho et al.[45] $\mu_{\theta}(x_{t},t)$ can be represented as:
$\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))$ (6)
$\beta_{\theta}(x_{t},t)$ can be represented as:
$\beta_{\theta}(x_{t},t)=\exp(\epsilon_{\theta}(x_{t},t)\log\beta_{t}+(1-\epsilon_{\theta}(x_{t},t))\log\tilde{\beta_{t}})$ (7)
Where $\tilde{\beta_{t}}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t}$. $\epsilon_{\theta}(x_{t},t)$is the value predicted by the network.
@@
According to Ho et al.[45], during single step training, only the $L_{t-1}$ loss needs to be computed. Hence, a simplified loss function can be derived as shown in Equation (9), where $\epsilon_{t}$ is given by Equation (4).
@@
Figure 2: Overall framework of the noise matching network.
@@ II-B Reverse sampling generation process:
Using equations (6) and (7), the reparameterization technique can be employed to obtain $x_{t-1}$.
$x_{t-1}$ $=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t))$ (11)
$+\epsilon\sqrt{\exp(\epsilon_{\theta}(x_{t},t)\log\beta_{t}+(1-\epsilon_{\theta}(x_{t},t))\log\tilde{\beta_{t}})},\epsilon\sim\mathcal{N}(0,\mathbf{I})$
@@
Figure 3: Backbone network architecture.
@@ III-A Constraint correction:
$\mu_{\theta}(x_{t},y,t)$ is represented as:
$\mu_{\theta}(x_{t},y,t)=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},y,t))$ (13)
$\beta_{\theta}(x_{t},y,t)$ is represented as:
$\beta_{\theta}(x_{t},y,t)=\exp(\epsilon_{\theta}(x_{t},y,t)\log\beta_{t}+(1-\epsilon_{\theta}(x_{t},y,t))\log\tilde{\beta_{t}})$ (14)
Where $\tilde{\beta_{t}}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t}$. The network is modified to estimate $\epsilon_{\theta}(x_{t},y,t)$. The value of $y$ is the self-supervised constraint we introduce into the constrained diffusion model.
@@
In the reverse sampling generation process, after introducing self-supervised constraints, the sampling at step t is:
Using equations (13) and (14), we can obtain $x_{t-1}$ using the reparameterization technique:
$x_{t-1}$ $=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},y,t))$ (18)
$+\epsilon\sqrt{\exp(\epsilon_{\theta}(x_{t},y,t)\log\beta_{t}+(1-\epsilon_{\theta}(x_{t},y,t))\log\tilde{\beta_{t}})},\epsilon\sim\mathcal{N}(0,\mathbf{I})$
Figure 4: Guided sampling.
@@ IV-A Evaluation Metrics:
Figure 7: (a) Ground truth,(b) 50% random missing, (b-1) to (b-4) are the reconstruction results of Unet, diffusion model, MDA GAN, and SeisFusion, respectively. (b-5) to (b-8) are the residuals of each reconstruction result relative to the ground truth.(c) 80% random missing, (c-1) to (c-4) are the reconstruction results of Unet, diffusion model, MDA GAN, and SeisFusion, respectively. (c-5) to (c-8) are the residuals of each reconstruction result relative to the ground truth.
@@ IV-C Ablation:
Figure 8: (a) Ground truth,(b) 50 continuous missing traces, (b-1) to (b-4) are the reconstruction results of Unet, diffusion model, MDA GAN, and SeisFusion, respectively. (b-5) to (b-8) are the residuals of each reconstruction result relative to the ground truth.(c) 100 continuous missing traces, (c-1) to (c-4) are the reconstruction results of Unet, diffusion model, MDA GAN, and SeisFusion, respectively. (c-5) to (c-8) are the residuals of each reconstruction result relative to the ground truth.
@@ IV-D Data Reconstruction:
During the sampling generation process, in Algorithm 1, T is set to 250 based on the state-of-the-art diffusion model, and u is set to 10. To validate the effectiveness of our proposed method, we selected three other models for comparative testing: Unet, diffusion model, and MDA GAN. Unet is the most commonly used model, which has been proven to have excellent performance in seismic data reconstruction[23]. The diffusion model[45] is directly extended from the 2D model, with training parameters consistent with our model. For MDA GAN[16], the authors provided the source code and weight files, stating that they did not need to be retrained. Therefore, we directly used the code and weights provided by the authors for reconstruction without any modifications.
@@ IV-D1 SEG C3
Random discrete missing: To simulate possible complex situations, we set two different missing rates, 50% and 80%, by setting traces to 0 to represent missing data. Fig. 7 shows the reconstruction results of different methods under these two missing rates. It can be observed that the reconstruction results of Unet already show low consistency at a 50% missing rate, and the discontinuity becomes more pronounced when the missing rate increases to 80%. Residual plots indicate a significant deviation between its reconstruction results and ground truth. The original diffusion model, lacking constraints and subject to sampling uncertainty, yields highly random and deviant reconstruction results compared to the ground truth. Both MDA GAN and our method perform well in reconstructing data under both missing scenarios, with a slight advantage of our method over MDA GAN, as indicated by residual plots.
@@
Random Missing 50% Traces UNet 1.4964e-4 34.1305 0.9613
Diffusion Model 1.3898e-3 25.3783 0.8751
@@
Random Missing 80% Traces UNet 1.7086e-3 30.7259 0.9326
Diffusion Model 5.1474e-3 22.3783 0.8595
MDA GAN 2.9637 35.2825 0.9789
SeisFusion 2.27954-4 36.4215 0.9881
Figure 9: (a) Ground truth,(b) 50% random missing, (b-1) to (b-4) are the reconstruction results of Unet, diffusion model, MDA GAN, and SeisFusion, respectively. (b-5) to (b-8) are the residuals of each reconstruction result relative to the ground truth.(c) 80% random missing, (c-1) to (c-4) are the reconstruction results of Unet, diffusion model, MDA GAN, and SeisFusion, respectively. (c-5) to (c-8) are the residuals of each reconstruction result relative to the ground truth.
Continuous missing: To evaluate the reconstruction performance of the proposed method under continuous missing, we set two scenarios of continuous missing, 50 continuous missing traces and 100 continuous missing traces. Traces were set to 0 to represent missing data. The reconstruction results of the four models are shown in Fig. 8. It can be observed that UNet fails to reconstruct seismic data under continuous missing, showing poor reconstruction performance. The original diffusion model also fails to complete the reconstruction task well due to the lack of constraints and sampling guidance under continuous missing. According to the residual maps, significant deviations from the ground truth can be observed in some areas. MDA GAN achieves relatively good reconstruction results, but deviations and inconsistencies still appear at the missing boundaries. Our method consistently provides the best reconstruction results under different missing scenarios.
Figure 10: (a) Ground truth,(b) 50 continuous missing traces, (b-1) to (b-4) are the reconstruction results of Unet, diffusion model, MDA GAN, and SeisFusion, respectively. (b-5) to (b-8) are the residuals of each reconstruction result relative to the ground truth.(c) 100 continuous missing traces, (c-1) to (c-4) are the reconstruction results of Unet, diffusion model, MDA GAN, and SeisFusion, respectively. (c-5) to (c-8) are the residuals of each reconstruction result relative to the ground truth..
@@
50 continuous missing traces UNet 0.0567 12.4661 0.7579
Diffusion Model 5.4142e-3 22.6646 0.8432
@@
100 continuous missing traces UNet 0.1116 9.5230 0.7095
Diffusion Model 0.0155 18.0705 0.6844

### v2

@@ Abstract
Seismic data often suffer from missing traces, and traditional reconstruction methods are cumbersome in parameterization and struggle to handle large-scale missing data. While deep learning has shown powerful reconstruction capabilities, convolutional neural networks’ point-to-point reconstruction may not fully cover the distribution of the entire dataset and may suffer performance degradation under complex missing patterns. In response to this challenge, we propose a novel diffusion model reconstruction framework tailored for 3D seismic data. To facilitate three-dimensional seismic data reconstruction using diffusion models, we introduce conditional constraints into the diffusion model, constraining the generated data of the diffusion model based on the input data to be reconstructed. We introduce a 3D neural network architecture into the diffusion model and refine the diffusion model’s generation process by incorporating existing parts of the data into the generation process, resulting in reconstructions with higher consistency. Through ablation studies determining optimal parameter values, although the sampling time is longer, our method exhibits superior reconstruction accuracy when applied to both field datasets and synthetic datasets, effectively addressing a wide range of complex missing patterns. Our implementation is available at https://github.com/WAL-l/SeisFusion.
@@ I Introduction
Figure 1: The diffusion model workflow involves $q$ for the forward encoding training process, where $x_{0}$ is gradually encoded into $x_{T}$. $p$ represents the reverse sampling generation process, where $x_{0}$ is gradually sampled from $x_{T}$, with $x_{t}^{inputtraces}$ serving as constraints added during the sampling process.
Traditional methods, although theoretically capable of data interpolation and reconstruction[17], heavily depend on manual parameter selection and struggle with interpolating large continuous missing data[18]. In recent years, with the development of deep learning in recent years, the use of deep learning for seismic data processing has attracted a lot of attention [19] [20] [21], and neural networks with different structures have been proposed for the reconstruction of seismic data[22, 23]. In the realm of 2D reconstruction, Wang et al. [14] advocated for using CAE to interpolate missing seismic data, while Chai et al. [24] employed the UNet network for seismic data reconstruction. Abedi et al.[25] propose a new ensemble deep model and a customized self supervised training method for reconstructing seismic data with continuous missing traces. The proposed model consists of two U-net branches, each receiving training data from different data transformation modules. These transformation modules enhance underrepresented features and promote diversity among learners. Application to two benchmark synthetic datasets and two real datasets demonstrates that the accuracy of U-net is improved compared to traditional U-net methods. Inspired by the competitive results using the conditional diffusion probabilistic model(DPM) in the field of image superresolution, Liu and Ma [26] proposed a DPM-based seismic data interpolation method. The known parts of the data are used as constraints for the DPM to reconstruct the seismic data, demonstrating strong reconstruction performance. Deng et al. [27] introduced constraint modifications into the diffusion model and improved the sampling process by incorporating resampling techniques, achieving better results compared to classical convolution methods. Wang et al. [28] proposed a reconstruction method based on a classifier-guided conditional seismic denoising diffusion probabilistic model. This method employs classifier-guided techniques during diffusion model inference sampling, exhibiting very high reconstruction accuracy. However, since real seismic data is inherently 3D, focusing solely on inline slices for 2D reconstruction neglects crucial crossline information. One by one reconstruction of 2D slices can introduce discontinuities in stacked slices, particularly in datasets featuring large-scale or continuous missing data and other complex scenarios. Therefore, opting for 3D reconstruction is a more pragmatic approach. Chai et al.[29] utilized an end-to-end 3D UNet for data reconstruction, observing superior performance compared to 2D methods. Qian et al.[30] proposed a deep tensor autoencoder, introducing tensor backpropagation, which exhibited commendable performance in reconstructing irregular data. Abedi and Pardo[17] proposed a multi-directional neural network that combines 2D and 3D learning capabilities. Initially, two networks are trained to perform simpler 2D reconstructions in the horizontal and vertical directions. Then, these networks are used as two parallel branches of a single network to perform the reconstruction of 3D data. Tests on synthetic data and field datasets show that this method achieves relatively accurate results and avoids the discontinuities associated with 2D U-Net. As GAN[31] showcase potent generative capabilities, utilizing improved GAN networks for seismic data interpolation has gained traction. Dou et al.[16] introduced a multidimensional adversarial generative adversarial network (MDA GAN), incorporating three discriminators and refining the loss function, resulting in robust reconstruction performance. Yu et al.[32] advocated for a CWGAN for data reconstruction, with results demonstrating its superiority over 3D UNet.
@@
1. Introduced a conditional supervision constraint into the diffusion model, using the known parts of the data as constraints to guide the generated data of the diffusion model, thereby ensuring higher consistency with the known data.
@@
3. Improved the generation and reconstruction process of the model by guiding the diffusion model’s sampling process with the known parts of the data. This approach ensures that the sampling results include prior information from the known data, enabling the generation of reconstruction data with higher consistency.
@@ II-A Forward Encoding Training Process:
Where $\beta_{t}$ is obtained from the existing variance table, typically set to linearly increase from 0.0001 to 0.002.
@@
where $\alpha_{t}=1-\beta_{t},\overline{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$.
@@
According to Ho et al.[48], during single step training, only the $L_{t-1}$ loss needs to be computed. Hence, a simplified loss function can be derived as shown in Equation (5), where $\epsilon_{t}$ is is sampled from a Gaussian distribution and $\epsilon_{\theta}(x_{t},t)$ is predicted by network.
@@
Figure 2: The overall framework of the noise matching network is as follows. $\epsilon_{\theta}(x_{t},y,t)$ is composed of the sum of the output obtained from inputting $x_{t}$ into the network (on the left) and the output obtained from inputting $y_{t}$ into the network (on the right). $y_{t}$ represents known data obtained by adding noise through the forward encoding process.
@@ II-B Reverse sampling generation process:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)),\beta_{\theta}(x_{t},t))$ (6)
@@
Figure 3: Backbone network architecture. On the left is the overall network architecture. The seismic data input undergoes a three-dimensional convolution, followed by downsampling through DownBlock and DownSample. After processing by MiddleBlock, it undergoes upsampling through UpBlock and UpSample, then passes through a three-dimensional convolution to obtain the output. On the right, from top to bottom, are the structures of DownBlock, MiddleBlock, and UpBlock.
@@ III-A Constraint correction:
The value of $y$ is the self-supervised constraint we introduce into the constrained diffusion model.
@@
The network changes from predicting $\epsilon_{\theta}(x_{t},t)$ to predicting $\epsilon_{\theta}(x_{t},y,t)$. In the reverse sampling generation process, after introducing self-supervised constraints, the sampling at step t is:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{\alpha}_{t}}}\epsilon_{\theta}(x_{t},y,t)),\beta_{\theta}(x_{t},y,t))$ (10)
Figure 4: Guided sampling process. Guided sampling modifies the standard denoising process in order to condition on the given seismic data content. In each step, we sample the known region (top) from the input and the reconstruction part from the diffusion model output described in the Constraint Correction subsection(bottom).
@@ IV-A Evaluation Metrics:
Figure 7: (a) Ground truth,(b) 50% random missing, (b-1) to (b-4) are the reconstruction results of Unet, conditional DPM, MDA GAN, and SeisFusion, respectively. (b-5) to (b-8) are the residuals of each reconstruction result relative to the ground truth.(c) 80% random missing, (c-1) to (c-4) are the reconstruction results of Unet, conditional DPM, MDA GAN, and SeisFusion, respectively. (c-5) to (c-8) are the residuals of each reconstruction result relative to the ground truth.
@@ IV-C Ablation:
Figure 8: (a) Ground truth,(b) 50 continuous missing traces, (b-1) to (b-4) are the reconstruction results of Unet, conditional DPM, MDA GAN, and SeisFusion, respectively. (b-5) to (b-8) are the residuals of each reconstruction result relative to the ground truth.(c) 100 continuous missing traces, (c-1) to (c-4) are the reconstruction results of Unet, conditional DPM, MDA GAN, and SeisFusion, respectively. (c-5) to (c-8) are the residuals of each reconstruction result relative to the ground truth.
Figure 9: (a) Ground truth,(b) 50% random missing, (b-1) to (b-4) are the reconstruction results of Unet, conditional DPM, MDA GAN, and SeisFusion, respectively. (b-5) to (b-8) are the residuals of each reconstruction result relative to the ground truth.(c) 80% random missing, (c-1) to (c-4) are the reconstruction results of Unet, conditional DPM, MDA GAN, and SeisFusion, respectively. (c-5) to (c-8) are the residuals of each reconstruction result relative to the ground truth.
@@ IV-D Data Reconstruction:
During the sampling generation process, in Algorithm 1, T is set to 250 based on the state-of-the-art diffusion model, and $U$ is set to 10. To validate the effectiveness of our proposed method, we selected three other models for comparative testing: Unet, conditional DPM, and MDA GAN. Unet is the most commonly used model, which has been proven to have excellent performance in seismic data reconstruction. We redesigned the UNet following the methodology of Anet[23] and trained it using the same dataset. However, since the related studies did not release their code, we could not ascertain the exact hyperparameters. We trained the model to its optimal state according to the parameters set in this paper, using Adam as the optimization algorithm with a learning rate of 1e-4. The training was conducted for 300 epochs on two RTX3090Ti GPUs. The conditional DPM[26] is two-dimensional and cannot be used for three-dimensional reconstruction, so we made some modifications by changing their 2D architecture to a 3D architecture, enabling it to perform 3D reconstruction. For MDA GAN[16], the authors provided the source code and weight files, stating that they did not need to be retrained. Therefore, we directly used the code and weights provided by the authors for reconstruction without any modifications.
@@ IV-D1 SEG C3
Figure 10: (a) Ground truth,(b) 50 continuous missing traces, (b-1) to (b-4) are the reconstruction results of Unet, conditional DPM, MDA GAN, and SeisFusion, respectively. (b-5) to (b-8) are the residuals of each reconstruction result relative to the ground truth.(c) 100 continuous missing traces, (c-1) to (c-4) are the reconstruction results of Unet, conditional DPM, MDA GAN, and SeisFusion, respectively. (c-5) to (c-8) are the residuals of each reconstruction result relative to the ground truth.
Random discrete missing: To simulate possible complex situations, we set two different missing rates, 50% and 80%, by setting traces to 0 to represent missing data. Fig. 7 shows the reconstruction results of different methods under these two missing rates. It can be observed that the reconstruction results of Unet already show low consistency at a 50% missing rate, and the discontinuity becomes more pronounced when the missing rate increases to 80%. Residual plots indicate a significant deviation between its reconstruction results and ground truth. On the other hand, when reconstructing three-dimensional data, conditional DPM experiences a slight performance decrease due to the increased complexity of the data. Both MDA GAN and our method perform well in reconstructing data under both missing scenarios, with a slight advantage of our method over MDA GAN, as indicated by residual plots.
@@
Random Missing 50% Traces UNet 1.2044e-4 36.8362 0.9651
Conditional DPM 8.9571e-5 39.2495 0.9864
@@
Random Missing 80% Traces UNet 8.2702e-4 32.2894 0.9451
Conditional DPM 4.8769e-4 33.1184 0.9791
MDA GAN 2.9637e-4 35.2825 0.9789
SeisFusion 2.2795e-4 36.4215 0.9881
Continuous missing: To evaluate the reconstruction performance of the proposed method under continuous missing, we set two scenarios of continuous missing, 50 continuous missing traces and 100 continuous missing traces. Traces were set to 0 to represent missing data. The reconstruction results of the four models are shown in Fig. 8. It can be observed that UNet fails to reconstruct seismic data under continuous missing, showing poor reconstruction performance. When reconstructing such continuous missing data, conditional DPM exhibits a bias towards higher values compared to the original data. According to the residual maps, significant deviations from the ground truth can be observed in some areas. MDA GAN achieves relatively good reconstruction results, but deviations and inconsistencies still appear at the missing boundaries. Our method consistently provides the best reconstruction results under different missing scenarios.
@@
50 continuous missing traces UNet 1.7565e-3 27.5535 0.9328
Conditional DPM 5.1959e-4 32.1184 0.9613
@@
100 continuous missing traces UNet 4.0417e-3 23.9342 0.8414
Conditional DPM 8.7652e-4 25.4631 0.9084